# Epstein Surfaces, $W$-Volume, and the Osgood-Stowe Differential

### v1

@@ 5.2 A flow on metrics
If $\hat{g}$ is a conformal metric on a projective structure we can define scaled metrics by $\hat{g}_{t}=e^{2t}\hat{g}$. Then as we will see below $\hat{B}_{t}=e^{-2t}\hat{B}$ are the projective shape operators for $\hat{g}_{t}$. The formulas for the dual metrics are more involved. We will eventually see that the dual metrics are the immersions obtained by flowing the original immersion in the normal direction.
Given a bundle endomorphism $B\colon TS\to TS$ define endomorphisms
$A_{t},C_{t}\colon TS\to TS$
by
$A_{t}=\cosh t\cdot{\operatorname{Id}}+\sinh t\cdot B\quad\mbox{ and }\quad C_{t}=\sinh t\cdot{\operatorname{Id}}+\cosh t\cdot B.$
We see next that these endomorphisms determine the pairs $(g_{t},B_{t})$ dual to $(\hat{g}_{t},\hat{B}_{t})$.
@@
[…]
@@
$\Box$
We remark that at a given point $p\in S$ the metrics $g_{t}$ can be singular (not positive definite) in which case the shape operator $B_{t}$ will will not be defined at $p$. However, for any given $p$ this can happen at most two values of $t$.
@@
Proof: By Theorem 5.12 and Proposition 5.4 $(f_{t})^{*}g_{{{\mathbb{H}}^{3}}}=g_{t}=A_{t}^{*}g$ and the shape operator for $f_{t}$ is $B_{t}=A_{t}^{-1}C_{t}$. The formula for $\lambda_{i}(t)$ follows.
@@
The one-parameter family of metrics $g_{t}$ could be combined to form a 3-manifold with a (possibly singular) Riemannian metric. Given an interval $I$ we define the flow space ${\cal F}_{(g,B)}(I)$ to be the product $S\times I$ with metric $g_{t}\oplus dt^{2}$.
Theorem 5.6
The $(g,B)$ satisfy the Gauss-Codazzi equations metric then the flow space ${\cal F}_{(g,B)}(I)$ has a metric of constant curvature $-1$ wherever it is non-singular. Furthermore $B$ is the shape operator for the surface $S\times\{0\}$ in ${\cal F}_{(g,B)}(I)$.
Proof: We let $\overline{\nabla}$ and $\nabla$ and $\bar{R}$ and $R$ be the connections and curvature tensors for the metrics $\overline{g}=g_{t}\oplus dt^{2}$ on the product $S\times I$ and $g$ on $S$. Let $n=\partial_{t}$ be a unit vector field orthogonal to the product surfaces. A vector field $X$ on $S\times I$ is horizontal if it is tangent to the level surfaces and $[n,X]=0$.
Let $(\hat{g},\hat{B})$ be the pair dual to $(g,B)$. Then the pairs $(e^{2t}\hat{g},e^{-2t}\hat{B})$ satisfy the projective Gauss-Codazzi equations by Theorem 5.1 and Proposition 5.4. Theorem 5.1 then implies that $(g_{t},B_{t})$ satisfy the Gauss-Codazzi equation in ${{\mathbb{H}}^{3}}$. Note that this computation could be made without using the dual metrics.
Next we check that if $X$ is a horizontal vector field then $B_{t}X=\bar{\nabla}_{X}n$. It suffices to check this when $t=0$. For this we have that by the compatibility of the metric with the connection we also have
$ng(X,X)=2g\left(\bar{\nabla}_{n}X,X\right)=2g\left(\bar{\nabla}_{X}n,X\right).$
Here we are using that ¯$\bar{\nabla}_{n}X=\bar{\nabla}_{X}n$ since $\bar{\nabla}$ is torsion free and $[X,n]=0$. Next we note that $ng(X,X)$ is the time zero derivative of
$A_{t}^{*}g(X,X)=\cosh^{2}t\cdot g(X,X)+2\sinh t\cdot\cosh t\cdot g(BX,X)+\sinh^{2}t\cdot g(BX,BX)$
so $ng(X,X)=2g(BX,X)$. Therefore $g(BX,X)=g\left(\bar{\nabla}_{X}n,X\right)$ for all horizontal $X$. Since both $B$ and $X\mapsto\nabla_{X}n$ are symmetric it follows that $BX=\nabla_{X}n$ and $B$ is the shape operator for $S\times\{0\}$ in ${\cal F}_{(g,B)}(I)$.
We note that the metric $\bar{g}$ has constant sectional curvature $-1$ if and only if
$\bar{R}(X,Y,Z,W)=-\left(\bar{g}(X,W)\bar{g}(Y,Z)-\bar{g}(X,Z)\bar{g}(Y,W)\right).$ (5.1)
This equation will hold if and only if it holds when $\{X,Y,Z,W\}$ vary over an orthonormal frame. Now choose horizontal vector fields $E_{1}$ and $E_{2}$ such that they are an orthonormal frame for $S$ at a point $p$ and $E_{3}=n$. Then $\{E_{1},E_{2},E_{3}\}$ are are an orthonormal frame for the product at $(p,0)$. Let $R_{ijkl}=R(E_{i},E_{j},E_{k},E_{l})$. By (5.1) we need to show that $R_{ijji}=-R_{ijij}=-1$ when $i\neq j$ and $R_{ijkl}=0$ otherwise.
The Bianchi identities greatly reduce the number of terms we need to compute. In particular we only need to find $R_{1221}$, $R_{1331}$, $R_{2332}$, $R_{1332}$ and $R_{ij3k}$ with $i,j$ and $k$ either 1 or 2. All other terms can either be reduced to these terms or are zero by the Bianchi identities.
We first calculate $R_{1221}$. This is the sectional curvature of the plane tangent to $S$. By the Gauss equation (for immersed surfaces in a Riemannian 3-manifold) this is the difference of the Gaussian curvature of $S$ and $\det(B)$, the determinant of the shape operator. As $(g,B)$ satisfies the Gauss equation in ${{\mathbb{H}}^{3}}$ the Gaussian curvature of $S$ is $\det(B)-1$ and it follows that $R_{1221}=-1$.
Note that for any $X$ and $Y$ by the Bianchi identities $\bar{R}(X,Y)n$ is tangent to $S$ (since $\bar{R}(X,Y,n,n)=0$) so
$\bar{R}(X,Y)n=\bar{\nabla}_{X}\bar{\nabla}_{Y}n-\bar{\nabla}_{Y}\bar{\nabla}_{X}n-\bar{\nabla}_{[X,Y]}n=\nabla_{X}\bar{\nabla}_{Y}n-\nabla_{Y}\bar{\nabla}_{X}n-\nabla_{[X,Y]}n.$
If $X$ and $Y$ are horizontal then so is $[X,Y]$ and therefore $[[X,Y],n]=0$ and
$\nabla_{[X,Y]}n=\nabla_{n}{[X,Y]}=B([X,Y])$
so
$\bar{R}(X,Y)n=\nabla_{X}(BY)-\nabla_{Y}(BX)-B([X,Y])$
and this is zero since $(g,B)$ satisfies the Codazzi equation. Therefore $R_{ij3k}=0$ when the $i,j$ and $k$ are either 1 or 2. Next we remark that since $(g,B)$ satisfies the Codazzi equation in ${{\mathbb{H}}^{3}}$ we have $R_{ij3k}=0$ when the $i,j$ and $k$ are in the set $\{1,2\}$.
If $X$ is horizontal we have
$\bar{R}(X,n)n=-\bar{\nabla}_{n}B_{t}X$
since ¯$\bar{\nabla}_{n}n=0$ and $[X,n]=0$. Therefore $\bar{R}(X,n,n,Y)=-\bar{g}(\bar{\nabla}_{n}B_{t}X,Y)$. We calculate this last term assuming that $Y$ is also horizontal. We have
$n\bar{g}(B_{t}X,Y)=\bar{g}(\bar{\nabla}_{n}B_{t}X,Y)+\bar{g}(B_{t}X,\bar{\nabla}_{n}Y)=\bar{g}(\bar{\nabla}_{n}B_{t}X,Y)+\bar{g}(B_{t}X,B_{t}Y).$
As above this is also the time zero derivative of
$\bar{g}(B_{t}X,Y)$ $=$ $A_{t}^{*}g(B_{t}X,Y)=g(C_{t}X,A_{t}Y)$
$=$ $\sinh t\cdot\cosh t\cdot g(X,Y)+(\sinh^{2}t+\cosh^{2}t)g(BX,Y)$
$+\sinh t\cdot\cosh t\cdot g(BX,BY).$
Therefore $ng(B_{t}X,Y)=g(X,Y)+g(BX,BY)$ and it follows that $\bar{R}(X,n,n,Y)=-g(X,Y)$. Therefore $\bar{R}_{1331}=\bar{R}_{2332}=-1$ and $\bar{R}_{1332}=0$. $\Box$
We will find the following classical result of E. Cartan very useful:
Theorem 5.7 (E. Cartan)
Let $(M,g)$ be a Riemannian manifold of constant sectional curvature $-1$. Then every point in $M$ has a neighborhood isometric to a neighborhood in ${\mathbb{H}}^{n}$. Furthermore if $M$ is simply connected then there is a local isometry from $(M,g)$ to ${\mathbb{H}}^{n}$. If $M$ is also complete then $M$ is isometric to ${\mathbb{H}}^{n}$.
@@
Proof: For both the first and second bullets the restrictions on the eigenvalues of $B$ imply that $A_{t}$ (and hence $g_{t}$) is non-singular. We now show completeness.
@@
given by taking the last coordinate is $1$-Lipschitz which implies that the sequence $t_{i}$ is convergent and lies in some bounded interval $[-T,T]$. We also have that the projection
${\cal F}_{(g,B)}([-T,T])\to S$
to the first coordinate is $e^{T}$-Lipschitz to the $g$ metric on $S$. To see this let $\lambda_{i}$ be the eigenvalues of $B$ then the eigenvalues of $A_{t}$ are $\lambda_{i}(t)=\cosh t+\lambda_{i}\sinh t$ and the Lipschitz constant is bounded by the reciprocal of the minimum of the eigenvalues for $t\in[-T,T]$. As $\lambda_{i}\in[-1,1]$ we have
@@
In particular, $p_{i}$ is convergent in $S$ so $(p_{i},t_{i})$ is convergent in ${\cal F}_{(g,B)}(I)$.
If the eigenvalues are in the interval $[-1,1]$ and $S$ is simply connected then ${\cal F}_{(g,B)}(\mathbb{R})$ is a complete, simply connected 3-manifold of constant curvature $-1$ so by Theorem 5.7 ${\cal F}_{(g,B)}(\mathbb{R})$ is isometric to ${{\mathbb{H}}^{3}}$. $\Box$
5.3 The normal flow
Let $(M,g)$ be a Riemannian manifold and $S$ a smooth, oriented manifold with $\dim S+1=\dim M$ and let $f\colon S\to M$ be an immersion. Then we let $\hat{f}\colon S\to T^{1}M\subset TM$ be the lift of $f$ to the unit tangent bundle where $\hat{f}(p)$ is the unit normal to the immersion at $f(p)$ (where the direction of the normal is given by the orientation). If ${\mathfrak{g}}_{t}\colon T^{1}M\to T^{1}M$ is the geodesic flow we can compose maps to a family $\hat{f}_{t}={\mathfrak{g}}_{t}\circ\hat{f}$. We can then compose again with the projection $\pi\colon T^{1}M\to M$ to get smooth maps $f_{t}=\pi\circ\hat{f}_{t}$. Then the $f_{t}$ are the normal flow of the immersion $f$. We would like to understand how the metric and shape operator evolve under this flow when $M$ is a hyperbolic 3-manifold.
We begin with a preliminary observation about the derivative of a vector field:
Lemma 5.9
Let $(M,g)$ be a Riemannian manifold with Riemannian connection normal-$\nabla$. If $X\colon M\to TM$ is a smooth vector field then the derivative $X_{*}\colon TM\to T(TM)$ is determined by $\nabla X$.
Proof: Fix $p\in M$. And let $E_{i}$ be a basis of $T_{p}M$ extended to vector fields on a neighborhood $U$ of $p$ with $\nabla_{Y}E_{j}=0$ at $p$ for all $Y\in T_{p}M$. The $E_{i}$ give a local trivialization giving coordinates for $TU\subset TM$. Then for $q\in U$ we have $X(q)=\sum f_{i}(q)E_{i}$ for smooth functions $f_{i}$. Then as a map from $U$ to $TU$ we have
$X(q)=(q,f_{1}(q),\dots,f_{n}(q)).$
For $Y\in T_{p}M$ we then have
$X_{*}(Y)=\left(Y,\sum Yf_{i}\cdot E_{i}\right).$
We also have
$\nabla_{Y}X=\sum Yf_{i}\cdot E_{i}+f_{i}\nabla_{Y}E_{i}=\sum Yf_{i}\cdot E_{i}$
and therefore
$X_{*}(Y)=\left(Y,\nabla_{Y}X\right)$
so $X_{*}$ only depends on $\nabla X$. $\Box$
The following is an immediate corollary:
Corollary 5.10
The derivative of $\hat{f}$ at $p$ is determined by the derivative $f_{*}$ and shape operator $B$ at $p$.
We now restrict to the case when $M={{\mathbb{H}}^{3}}$ and $S$ is a surface. One consequence of this computation is that an immersed surface is determined by its first and second fundamental forms.
Theorem 5.11
Let $(g,B)$ be a metric and tangent bundle endomorphism on a connected surface $S$ that satisfy the Gauss-Codazzi equations in ${{\mathbb{H}}^{3}}$. Let
$f_{0},f_{1}\colon S\to{{\mathbb{H}}^{3}}$
be immersions such that $g=(f_{i})^{*}g_{{\mathbb{H}}^{3}}$ and $B$ is the shape operator for both immersions. The $f_{0}$ and $f_{1}$ differ by post-composition with an isometry of ${{\mathbb{H}}^{3}}$.
Proof: Let $(g^{i}_{t},B^{i}_{t})$ be the metrics and shape operators for the time $t$ flow of $f_{0}$ and $f_{1}$. We’ll first show that they are equal. It is enough to do this at a point $p\in S$. We normalize so that $f_{0}(p)=f_{1}(p)$. Since the metrics are equal we can further normalize so that the derivatives agree at $p$. Then $\hat{f}_{0}$ and $\hat{f}_{1}$ agree at $p$ and since the shape operators agree we also have by Corollary 5.10 that the derivatives of $\hat{f}_{0}$ and $\hat{f}_{1}$ agree at $p$. By the chain rule the derivatives of the time $t$ flows $\left(\hat{f}_{i}\right)_{t}$ are the composition of the derivative of ${\mathfrak{g}}_{t}$ and the derivative of $\hat{f}_{i}$. Therefore the derivatives of $\left(\hat{f}_{0}\right)_{t}$ and $\left(\hat{f}_{1}\right)_{t}$ agree at $p$. Since this derivative determines both the metric and shape operator we have that $(g^{0}_{t},B^{0}_{t})$ and $(g^{1}_{t},B^{1}_{t})$ agree at $p$ and, hence, everywhere on $S$.
Now define maps $F_{i}\colon S\times(-\epsilon,\epsilon)\to{{\mathbb{H}}^{3}}$ by $F_{i}(p,t)=(f_{i})_{t}(p)$. By the above paragraph $F^{*}_{0}g_{{\mathbb{H}}^{3}}=F^{*}_{1}g_{{\mathbb{H}}^{3}}$. We also have that the $F_{i}$ are immersions at $(p,0)$ for all $p\in S$. Therefore we can choose a neighborhood $U$ of $(p,0)$ in $S\times(-\epsilon,\epsilon)$ such that $F_{i}$ are embeddings on $U$. Therefore $F_{1}\circ F_{0}^{-1}$ is an isometry from $F_{0}(U)$ to $F_{1}(U)$. As the $F_{i}(U)$ are open sets in ${{\mathbb{H}}^{3}}$ and any isometry between open sets extends to an isometry of ${{\mathbb{H}}^{3}}$ we get an isometry $\phi_{p}$ of all of ${{\mathbb{H}}^{3}}$ with $\phi_{p}\circ F_{0}=F_{1}$ on $U$. However, $\phi_{p}$ will be constant as we vary $p$ so there will be single isometry $\phi$ with $\phi\circ F_{0}=F_{1}$ on an open neighborhood of $S\times\{0\}$. Restricting to $S\times\{0\}$ we have $\phi\circ f_{0}=f_{1}$ as claimed. $\Box$
We can now derive the classical result of Bonnet:
Corollary 5.12 (Bonnet)
$f\colon S\to{{\mathbb{H}}^{3}}$
Proof: Fix an interval $I$ that contains 00. Then for every $p\in S$ there is a neighborhood of $p\times\{0\}$ in ${\cal F}_{(g,B)}(I)$ where the metric is non-singular. Using this we can find a simply connected neighborhood $U$ of $S\times\{0\}$ in ${\cal F}_{(g,B)}(I)$ where the metric is non-singular. By Theorem 5.1 the metric on this neighborhood has constant sectional curvature equal to $-1$ and so by Theorem 5.7 there is local isometry from $U$ to ${{\mathbb{H}}^{3}}$. The restriction of this isometry to $S\times\{0\}$ gives the immersion of $S$ in ${{\mathbb{H}}^{3}}$ and the uniqueness comes from Theorem 5.11. $\Box$
@@ 5.4 The hyperbolic Gauss map
For an immersion $f$ the composition ${\mathfrak{g}}_{\infty}\circ\hat{f}$ gives a map from $S$ to $\widehat{{\mathbb{C}}}$.
We now consider the pushforward of $\hat{g}$ by the Gauss map $f_{\infty}$ and show that it is conformal on $\hat{\mathbb{C}}$ and can be described in terms of the visual metric.
For each $p\in{{\mathbb{H}}^{3}}$ the sphere $T^{1}_{p}{{\mathbb{H}}^{3}}$ has an invariant metric (under isometries of ${{\mathbb{H}}^{3}}$ that fix $p$) that is unique up to scale. This is a spherical metric and we fix the scale so that the sphere has radius $1$. The visual metric $\nu_{p}$ based at $p$ is the push-forward of this measure via the hyperbolic Gauss map. We observe that in the upper half space model if $p=(0,1)$ then
$\nu_{(0,1)}=\frac{4{g_{\rm euc}}}{(1+|z|^{2})^{2}}.$
@@
$f\colon S\to{{\mathbb{H}}^{3}}$ be an immersion with shape operator $B$ and let $\hat{f}$ be the lift to $T^{1}{{\mathbb{H}}^{3}}$. We also let $g=f^{*}g_{{{\mathbb{H}}^{3}}}$. If $f_{\infty}={\mathfrak{g}}_{\infty}\circ\hat{f}$ then $f_{\infty}$ is an immersion at $p\in S$ if and only if $-1$ is not an eigenvalue of $B$ at $p$. Furthermore at $p$ we have
@@
so $f_{\infty}$ is a conformal map from $\hat{g}$ to $-^\widehat{{\mathbb{C}}}$.
Proof: Fix a point $p\in S$ and choose $T>0$ such that the eigenvalues of $B_{T}$ at $p$ are $\geq-1$. The flow space ${\cal F}_{(g,B)}([T,\infty))$ is the manifold $S\times[T,\infty)$ with the warped product metric $g_{t}\oplus dt^{2}$ and there is a smooth map $F\colon{\cal F}_{(g,B)}([T,\infty))\to{{\mathbb{H}}^{3}}$ with $g_{t}\oplus dt^{2}=F^{*}g_{{{\mathbb{H}}^{3}}}$. From our choice of $T$ we have that $F$ is an immersion along the geodesic $\{p\}\times[T,\infty)$. Furthermore $F$ extends smoothly to a map from $S\times\{\infty\}$ to $\widehat{{\mathbb{C}}}$ by setting $F(q,\infty)=f_{\infty}(q)$ for all $q\in S$. The scaled metric $4e^{-2t}g_{t}\oplus dt^{2}$ then extends to continuously to $S\times\{\infty\}$ where the limiting metric is $\hat{g}=({\operatorname{Id}}+B)^{*}g$. Working in the upper half space model we then arrange the map $F$ such that $F(p,t)=f_{t}(p)=(0,e^{-t})$ and $f_{\infty}(p)=0$. Note that $g_{{{\mathbb{H}}^{3}}}=\frac{1}{s^{2}}g_{{\mathbb{E}}^{3}}$ (in the upper half space coordinates $(z,s)$ for ${{\mathbb{H}}^{3}}$). Therefore, given our normalization, along the geodesic $\{p\}\times[T,\infty)$ in ${\cal F}_{(g,B)}([T,\infty))$ we have that $4e^{-2t}g_{t}\oplus dt^{2}=4F^{*}g_{{\mathbb{E}}^{3}}$. It then follows by continuity that $p\in S$ and $0\in\mathbb{C}$ we have $\hat{g}=4(f_{\infty})^{*}{g_{\rm euc}}=(f_{\infty})^{*}\nu_{(0,1)}$. Since $f(p)=(0,1)$ we have $\nu_{(1,0)}=\nu_{f(p)}$ and the formula follows. Finally since $\hat{g}$ is singular at $p$ if and only if $-1$ is an eigenvalue of $B$ at $p$ we have that $f_{\infty}$ is an immersion at $p$ if and only if $-1$ is not an eigenvalue of $B$ at $p$. $\Box$

### v2

@@ 5.4 A flow on metrics
If $\hat{g}$ is a conformal metric on a projective structure we can define scaled metrics by $\hat{g}_{t}=e^{2t}\hat{g}$. Then as we will see below $\hat{B}_{t}=e^{-2t}\hat{B}$ are the projective shape operators for $\hat{g}_{t}$. The formulas for the dual metrics are more involved. We now see that the dual metrics are the normal flow immersions for the original immersion.
@@
$\hat{B}_{t}=e^{-2t}\hat{B},\qquad g_{t}=A_{t}^{*}g,\quad\mbox{and}\quad B_{t}=\left(A_{t}\right)^{-1}C_{t}.$
In particular if $f:S\rightarrow{{\mathbb{H}}^{3}}$ is an immersion with fundamental pair $(g,B)$ dual to $(\hat{g},\hat{B})$ then the normal flow map $f_{t}:S\rightarrow{{\mathbb{H}}^{3}}$ has fundamental pair $(g_{t},B_{t})$ dual to $(e^{2t}\hat{g},e^{-2t}\hat{B})$.
@@
The rest follows from Theorem 5.9. $\Box$
We remark that at a given point $p\in S$ the metrics $g_{t}$ can be singular (not positive definite) in which case the shape operator $B_{t}$ will not be defined at $p$. However, for any given $p$ this can happen at most two values of $t$.
@@
Proof: By Theorem 5.9 $(f_{t})^{*}g_{{{\mathbb{H}}^{3}}}=g_{t}=A_{t}^{*}g$ and the shape operator for $f_{t}$ is $B_{t}=A_{t}^{-1}C_{t}$. The formula for $\lambda_{i}(t)$ follows.
@@ 5.5 The flow space of a pair $(g,B)$
The one-parameter family of metrics $g_{t}$ can be combined to form a 3-manifold with a (possibly singular) Riemannian metric. We define the flow space ${\cal F}_{(g,B)}(I)$ to be the product $S\times I$ with metric $g_{t}\oplus dt^{2}$. We have the following corollary.
Corollary 5.12
Let $(g,B)$ satisfy the Gauss-Codazzi equations on $S$. Then ${\cal F}_{(g,B)}(I)$ is a hyperbolic manifold wherever the metric $g_{t}$ is non-singular. If $S$ is simply connected then there is a smooth map $F:S\times I\rightarrow{{\mathbb{H}}^{3}}$ such that $F^{*}g_{{{\mathbb{H}}^{3}}}=g_{t}\oplus dt^{2}$ and $F$ is unique up to post-composition by a Möbius transformation.
Proof: We first observe that for $S$ simply connected then by Theorem 5.9, we have the smooth map $F:S\times I\rightarrow{{\mathbb{H}}^{3}}$ defined by $F(q,t)=f_{t}(q)$ and $F^{*}g_{{{\mathbb{H}}^{3}}}=g_{t}\oplus dt^{2}$. By uniqueness of immersions, $F$ restricted to $S\times\{0\}$ is unique up to post-composition by Möbius transformation. As $F$ is defined by normal flow on this immersion, it follows that $F$ is unique up to post-composition by a Möbius tranformation. If $S$ is not simply connected then we obtain a hyperbolic structure on the non-singular points of ${\cal F}_{(g,B)}(I)$ by choosing for each $U\subseteq S$ open simply connected the map $F_{U}:U\times I\rightarrow{{\mathbb{H}}^{3}}$. By uniqueness up to post-composition, this gives an atlas for a hyperbolic structure on the non-singular set. $\Box$
@@
Proof: For both the first and second bullets the restrictions on the eigenvalues of $B$ imply that $A_{t}$ (and hence $g_{t}$) is non-singular. Thus the flow space is a hyperbolic manifold. We now show completeness.
@@
given by taking the last coordinate is $1$-Lipschitz which implies that the sequence $t_{i}$ is convergent and lies in some bounded interval $I\cap[-T,T]$. We also have that the projection
${\cal F}_{(g,B)}(I\cap[-T,T])\to S$
to the first coordinate is $e^{T}$-Lipschitz to the $g$ metric on $S$. To see this let $\lambda_{i}$ be the eigenvalues of $B$ then the eigenvalues of $A_{t}$ are $\lambda_{i}(t)=\cosh t+\lambda_{i}\sinh t$ and the Lipschitz constant is bounded by the reciprocal of the minimum of the eigenvalues for $t\in I\cap[-T,T]$. For $I=[0,\infty)$ we have $\lambda_{i}\geq-1$ and for $I=\mathbb{R}$, $\lambda_{i}\in[-1,1]$. In both cases we have for $t\in I\cap[-T,T]$
@@
In particular, $p_{i}$ is convergent in $S$ so $(p_{i},t_{i})$ is convergent in ${\cal F}_{(g,B)}(I)$. Thus in both cases, ${\cal F}_{(g,B)}(I)$ is complete.
Finally if $S$ is simply connected and the eigenvalues are in the interval $[-1,1]$ then ${\cal F}_{(g,B)}(\mathbb{R})$ is a simply connected, complete hyperbolic manifold. Therefore by Corollary 5.12. we have a local isometry $F:S\times\mathbb{R}\rightarrow{{\mathbb{H}}^{3}}$. As a local isometry between complete hyperbolic manifolds is a covering map, then as $S\times\mathbb{R}$ is simply connected, $F$ must be an isometry. $\Box$
@@ 5.6 The hyperbolic Gauss map
The map ${\mathfrak{g}}_{\infty}$ is particularly simple in the Minkowski model. We let ${\cal L}\subset\mathbb{R}^{3,1}$ be the light cone
=${\cal L}=\{x\in\mathbb{R}^{3,1}\ |\ {\langle}x,x{\rangle}=0,\ x\neq 0\}.$
If ${{\mathbb{H}}^{3}}\subseteq\mathbb{R}^{3,1}$ is the hyperboloid, then the Klein model is $P({{\mathbb{H}}^{3}})\subseteq\mathbb{RP}^{3}$, the projective hyperboloid with boundary $P({\cal L})$, the projective light cone. Then by the equation for geodesic flow in ${{\mathbb{H}}^{3}}$ (see equation 5.3)
${\mathfrak{g}}_{\infty}(q,v)=\lim_{t\rightarrow\infty}[\cosh(t)q+\sinh(t)v]=[q+v].$
We therefore define $\hat{\mathfrak{g}}_{\infty}:T_{1}{{\mathbb{H}}^{3}}\rightarrow{\cal L}$ by
$\hat{\mathfrak{g}}_{\infty}(q,v)=q+v.$
The identification $\partial{{\mathbb{H}}^{3}}=\widehat{{\mathbb{C}}}$ is then given by the map $p\colon{\cal L}\to\mathbb{S}^{2}\subset\mathbb{R}^{3}$
$p(x_{1},x_{2},x_{3},x_{4})=(x_{1}/x_{4},x_{2}/x_{4},x_{3}/x_{4})$
which descends to a homeomorphism on $P({\cal L})$. Therefore $^{\mathfrak{g}}_{\infty}:T_{1}{{\mathbb{H}}^{3}}\rightarrow\widehat{{\mathbb{C}}}$ satisfies
${\mathfrak{g}}_{\infty}=p\circ\hat{\mathfrak{g}}_{\infty}.$
We will need the following lemma.
@@
If $v,w\in T_{(x_{1},x_{2},x_{3},x_{4})}{\cal L}$ then
${\langle}v,w{\rangle}=\frac{1}{x_{4}^{2}}{\langle}p_{*}v,p_{*}w{\rangle}_{\rm{euc}}.$
Proof: For $u=(u_{1},u_{2},u_{3},u_{4})\in\mathbb{R}^{3,1}$ let $\hat{u}=(u_{1},u_{2},u_{3})\in\mathbb{R}^{3}$. Thus if $u\in T_{x}{\cal L}$
$p_{*}(u)=\frac{1}{x_{4}}\left(\hat{u}-\frac{u_{4}}{x_{4}}\hat{x}\right).$
As $v,w\in T_{x}{\cal L}$ then ${\langle}x,x{\rangle}={\langle}v,x{\rangle}={\langle}w,x{\rangle}=0$ giving
${\langle}\hat{x},\hat{x}{\rangle}_{\rm{euc}}=x^{2}_{4},\qquad{\langle}\hat{v},\hat{x}{\rangle}_{\rm{euc}}=v_{4}x_{4},\qquad{\langle}\hat{w},\hat{x}{\rangle}_{\rm{euc}}=w_{4}x_{4}.$
Thus
${\langle}p_{*}v,p_{*}w{\rangle}_{\rm{euc}}$ $=\frac{1}{x_{4}^{2}}\left\langle\hat{v}-\frac{v_{4}}{x_{4}}\hat{x},\hat{w}-\frac{w_{4}}{x_{4}}\hat{x}\right\rangle_{\rm{euc}}$
$=\frac{1}{x_{4}^{2}}\left({\langle}\hat{v},\hat{w}{\rangle}_{\rm{euc}}-\frac{w_{4}}{x_{4}}{\langle}\hat{v},\hat{x}{\rangle}_{\rm{euc}}-\frac{v_{4}}{x_{4}}{\langle}\hat{w},\hat{x}{\rangle}_{\rm{euc}}+\frac{v_{4}w_{4}}{x_{4}^{2}}{\langle}\hat{x},\hat{x}{\rangle}_{\rm{euc}}\right)$
$=\frac{1}{x_{4}^{2}}\left({\langle}\hat{v},\hat{w}{\rangle}_{\rm{euc}}-v_{4}w_{4}\right)=\frac{1}{x^{2}_{4}}{\langle}v,w{\rangle}$
$\Box$
For an immersion $f$ the composition $f_{\infty}={\mathfrak{g}}_{\infty}\circ\hat{f}$ gives a map from $S$ to $\widehat{{\mathbb{C}}}$ called the Gauss map of $f$. We now consider the pushforward of $\hat{g}$ by the Gauss map $f_{\infty}$ and show that it is conformal on $\hat{\mathbb{C}}$ and can be described in terms of the visual metric.
For each $p\in{{\mathbb{H}}^{3}}$ the sphere $T^{1}_{p}{{\mathbb{H}}^{3}}$ has an invariant metric (under isometries of ${{\mathbb{H}}^{3}}$ that fix $p$) that is unique up to scale. This is a spherical metric and we fix the scale so that the sphere has radius $1$. The visual metric $\nu_{p}$ based at $p$ is the push-forward of this measure via the hyperbolic Gauss map.
Lemma 5.16
Let $f\colon S\to{{\mathbb{H}}^{3}}$ be an immersion with shape operator $B$ and let $\hat{f}$ be the lift to $T^{1}{{\mathbb{H}}^{3}}$. We also let $g=f^{*}g_{{{\mathbb{H}}^{3}}}$. If $f_{\infty}={\mathfrak{g}}_{\infty}\circ\hat{f}$ then $f_{\infty}$ is an immersion at $p\in S$ if and only if $-1$ is not an eigenvalue of $B$ at $p$. Furthermore at $p$ we have
@@
so $f_{\infty}$ is a conformal map from $\hat{g}$ to $\widehat{{\mathbb{C}}}$.
Proof: By definition of $f_{\infty}$ we have
$f_{\infty}={\mathfrak{g}}_{\infty}\circ\hat{f}=p\circ\hat{\mathfrak{g}}_{\infty}\circ\hat{f}.$
Let $\hat{f}:S\rightarrow T\mathbb{R}^{3,1}$ be given by $\hat{f}(q)=(f(q),n(q))$. Then
$(\hat{\mathfrak{g}}_{\infty})_{*}(v,w)=v+w\qquad(\hat{f})_{*}(v)=(f_{*}(v),f_{*}(Bv))$
and therefore
$(\hat{\mathfrak{g}}_{\infty}\circ\hat{f})_{*}(v)=f_{*}(v)+f_{*}(Bv)=f_{*}(({\operatorname{Id}}+B)v).$
By post-composition by an isometry, we can assume that $f(q)=(0,0,0,1)$. Then $\nu_{f(q)}$ is the restriction of the euclidean metric to $\mathbb{S}^{2}$. Therefore
$\hat{g}(v,w)$ $=g(({\operatorname{Id}}+B)v,({\operatorname{Id}}+B)w)$
$={\langle}f_{*}(({\operatorname{Id}}+B)v),f_{*}(({\operatorname{Id}}+B)w){\rangle}$
$={\langle}p_{*}(f_{*}(({\operatorname{Id}}+B)v)),p_{*}(f_{*}(({\operatorname{Id}}+B)w)){\rangle}_{\rm{euc}}$
$={\langle}(f_{\infty})_{*}v,(f_{\infty})_{*}w{\rangle}_{\rm{euc}}$
$=(f_{\infty})^{*}(\nu_{f(q)})(v,w)$
$\Box$